OpenStudy (anonymous):

What is the value of the expression |a + b| + |c| when a = –5, b = 8, and c = –13? A. –26 B. 0 C. 16 D. –10

OpenStudy (anonymous):

\[\left| 8-5 \right| + \left| 13 \right|=3 +13 = 16\] c.